OpenStudy (anonymous):

Area of a circle with a radius of 28cm?

hartnn (hartnn):

use the formula \(\Large A_{circle} = \pi r^2 \) where r = radius = 28

OpenStudy (anonymous):

Okay, thanks! =3

hartnn (hartnn):

welcome ^_^ so what answer did u get ? :)

OpenStudy (anonymous):

2463, but I'm checking my work now o3o I'm not the best with math,

hartnn (hartnn):

it is correct :) use 2463 cm^2

OpenStudy (anonymous):

Ah, okay.

OpenStudy (anonymous):

A= \[A=\pi r^2\]

OpenStudy (anonymous):

A=3.14x28^2 =2461.76cm^2

OpenStudy (nevermind_justschool):

Don't forget to close the question :) Thanks :)